# AI Makes You Smarter, But None The Wiser: The Disconnect Between Performance and Metacognition

### v1

@@ 3.2. Materials
Participants’ logical reasoning ability was measured with the 20 multiple-choice logical reasoning problems used by Jansen et al. (2021) to approximate the Law School Admission Test (LSAT), a widely recognized, real-world assessment used in high-stakes decision-making, such as law school admissions (Shultz and Zedeck, 2011; Wainer, 1995). It also serves as a benchmark in machine learning research, making it ideal for comparing AI-assisted performance (Katz et al., 2024). Using the same items enabled us to compare our results to a representative sample of participants who had no AI-assistance in the task and replicate the results of the original study by Kruger and Dunning (1999).
@@
3.3. Task
Participants completed 20 LSAT items in a randomized order, with the problem displayed on the left-hand side of the screen and a ChatGPT interface on the right. For each problem, participants were asked to rate their confidence in their response (“How confident are you that your response is correct?”; from “unsure” to “certain” on a 100-step slider). To ensure AI use, participants were required to prompt ChatGPT at least once, wait for its response, answer the problem and indicate their confidence. Unlimited text chat interaction was enabled during the task.
3.4. Apparatus
@@
We inspected the software and flow of Jansen et al. (2021) and carefully replicated it, integrating a side-by-side view of ChatGPT and survey interface (Figure 2). We used ChatGPT-4o (gpt-4o-2024-05-13) due to its widespread use in enhancing cognitive task performance (Bastani et al., 2024; Draxler et al., 2024, 2023). A custom interface (Figure 2) was built to log user interactions, enabling qualitative chat analysis. The application automatically collected survey responses and chat logs and recorded them at the end of the study. We included a button to copy each problem and its answer options to the clipboard so they could be easily pasted in the chat.
3.5. Procedure
Upon entering the study on Prolific, participants were redirected to our application. After consenting to participate, they were asked to complete the expectations questionnaire. Then, they were briefly introduced to the task and allowed to test the chat interaction.
After solving the problems in a randomized order (3.3), participants were again asked to complete the expectations questionnaire, this time in the past tense. They also responded to the SNAIL questionnaire (Laupichler et al., 2023) and filled in their demographic information, including age, gender, occupation, education level, English proficiency, and whether they had taken the LSAT before. The study took, on average, 42 minutes to complete.
4. Results
@@
We analyzed the data in five steps. First, we compared our sample performance to that of Jansen et al. (2021). Second, we analyzed metacognitive accuracy – the difference between objective and estimated performance – and confidence ratings to assess metacognitive sensitivity. We then correlated metacognition metrics with performance and AI literacy measures. Fourth, we used a computational model to compare our sample and Jansen et al. (2021) to estimate how AI affects the DKE. Lastly, we qualitatively analyzed participant strategies  [^2] Given the large samples in our study and the ceiling effects encountered 5(a), we do not test for the normality of residuals of our variables. Instead, we model the data using a final Bayesian computational framework, which allows for more flexible assumptions and can account for ceiling effects in our main analysis. This approach provides more robust estimates by incorporating uncertainty in a probabilistic manner, rather than relying on strict parametric assumptions..
@@
Note: M (SD) and the sample size (n). Scale for the assessment of non-experts’ AI literacy (SNAIL). According to task context, rank % instructions are scaled as follows: marking 90% means you perform better than 90% of participants, marking 10% means you perform better than only 10% of participants, and marking 50% means you will perform better than half of the participants.
4.1. Human-AI Composite Performance
Figure 3. Comparison of performance scores between participants interacting with ChatGPT and a dataset without AI-assistance (Jansen et al., 2021). The blue density curve represents the distribution of performance scores in a sample of 246 participants using ChatGPT, showing a peak in performance at around 14 points. The yellow density curve corresponds to a larger sample (n = 3543) from the Jansen et al. (2021) dataset, with lower overall performance scores. The vertical dashed line indicates the mean performance score in the ChatGPT simulation.
@@ 4.1. Human-AI Composite Performance
The graph compares performance scores between participants using ChatGPT and those without AI assistance from the Jansen et al. dataset. The ChatGPT group shows higher performance, peaking around 14 points, while the no-AI group has lower overall scores. The dashed line marks the mean performance in the ChatGPT simulation.
@@
To see whether there is a synergy effect of using ChatGPT in the LSAT, we compared the average ChatGPT performance (100 runs at $M$ = 13.65;) to our user’s average performance ($M$ = 12.98, $SD$ = 2.88). We find a significant difference, with participants performing slightly worse than ChatGPT alone $t(245)=-3.66,p<.001,d=-0.23$ in the task. Next, we compared our sample’s performance to Jansen et al. (2021) representative sample of 3543 participants, who completed the same task without any assistance ($M$ = 9.45, $SD$ = 3.59). We find that in our sample, participants performed significantly better with ChatGPT assistance $t(245)=19.23,p<.001,d=1.23$ as compared to the Jansen et al. (2021) sample.
@@
Note that the difference between our sample’s and ChatGPT’s performance is rather small, at less than one point. 55.28% (136 of 246) of our participants performed better than ChatGPT. However, 89.43% (220 of 246) in our sample performed better than the average score of the Jansen et al. (2021) sample (see also Figure 3).
@@
Therefore, while overall performance increased with the use of ChatGPT, the composite performance of ChatGPT and the participant overtook the performance of ChatGPT alone for slightly more than half of the participants in our sample. With our prototype’s ability to enhance performance established, we can now turn our focus to investigating metacognitive abilities.
4.2. Metacognitive accuracy and sensitivity
@@
Participants were inaccurate in assessing their performance after task completion. On average, they estimated solving about 17 out of 20 items ($M$ = 16.50, $SD$ = 3.72). This overestimation of about 4 points could be distinguished from 0, $t(256)=14.98,p<.001,d=0.93$. Mean confidence (rated on a scale from 0-100) for correct answers was 82.49 (14.24) and for incorrect answers 77.00 (16.52), t(244) = 8.21, p ¡ .001, d = 0.52. To evaluate sensitivity, we conducted a Receiver Operating Characteristic (ROC) analysis. ROC analysis is a technique used to assess the performance of a judgment by plotting the true positive rate against the false positive rate across different thresholds. Here, we applied ROC analysis to understand how well participants’ confidence scores predicted whether their responses were correct (see Figure 4A). The effectiveness of the confidence scores for each participant in predicting correctness can be summarized as the Area Under the Curve (AUC). A higher AUC value indicates that the confidence scores are more reliable in reflecting the participants’ correctness. While most participants’ (210 out of 246; 85.37%) metacognitive AUC values are above .50 (random guessing), some participants’ confidence ratings were not able to distinguish between correct or incorrect trials (for the distribution of AUC values, refer to Figure 4B).
@@
Two plots related to metacognitive sensitivity. Plot A displays Receiver Operating Characteristic (ROC) curves for participants, illustrating the relationship between True Positive Rate (Sensitivity) and False Positive Rate. The dashed diagonal line represents random guessing. For each participant, we generated a ROC curve, as well as one from pooled responses, which illustrates the trade-off between the true positive rate. Plot B shows the distribution of Area Under the Curve (AUC) values across all participants, with a peak around 0.6, indicating variability in metacognitive sensitivity. Most participants perform slightly above chance level (AUC = 0.5).
4.3. Correlation of metacognitive ability, performance, and AI literacy (SNAIL)
A number of significant relationships were found when correlating several metacognitive indices with LSAT performance and AI literacy. We observed a strong positive relation between performance and participants’ average confidence estimates. Participants who perform well are also more confident on average. However, those who are, on average, more confident also overestimate their performance due to increased metacognitive bias. This is probably due to the relationship between SNAIL factors and performance estimates, where participants who express more technical knowledge and more critical appraisal also estimate their performance to be relatively higher. However, those with high technical understanding are also less accurate in their metacognitive judgements. All SNAIL factors correlated positively to average confidence. Note that these correlations are rather small and should thus be interpreted with caution. Metacognitive sensitivity (AUC and $\Delta conf$) was not related to AI literacy, performance or metacognitive accuracy.
@@ 4.3. Correlation of metacognitive ability, performance, and AI literacy (SNAIL)
$\Delta conf$ 0.04 0.03 -0.01
$\mu conf$ 0.24*** 0.46*** 0.27*** 0.10
AUC 0.03 0.05 0.03 -0.59*** -0.08
SNAIL TU 0.21** 0.17** -0.06 0.12 0.13* -0.10
SNAIL CA 0.10 0.14* 0.06 -0.04 0.25*** 0.03 0.49***
SNAIL PA 0.05 0.10 0.06 -0.01 0.24*** 0.05 0.57*** 0.81***
@@
4.4. AI use cancels the Dunning-Kruger effect
@@
We calculated the difference between quantiles to test whether metacognitive accuracy was worse in the low-scoring quantile than in the high-scoring quantile. Both quantile’s metacognitive accuracy differed from 0, (Q1: $t(109)=-10.15,p<.001,d=-0.97$, Q4: $t(21)=-3.64,p=.002,d=-0.78$), probably due to the overall bias. However, we found that the difference for Q1 is larger than for Q4, $t(130)=2.79,p=.006,d=0.49$ (see also 5(b) and Table 1). Note that this pattern of effect could be driven by metacognitive bias alone. To establish a DKE, we must first quantify the metacognitive noise in our sample. To do so, we employ a Bayesian computational model  [^4] For a guide on Bayesian techniques, see (Schad et al., 2021; van de Schoot et al., 2021; Dix, 2022; Bürkner, 2017; Kay et al., 2016), we used the tutorial of Nathaniel Haine’s as a starting point for our modelling efforts: http://haines-lab.com/post/2021-01-10-modeling-classic-effects-dunning-kruger/.
We developed a hierarchical Bayesian model to jointly estimate participants’ objective and perceived performance while accounting for latent skill, metacognitive bias, and metacognitive noise  [^5] For a theoretical model, see Burson et al. (2006). To allow for a baseline comparison of the DKE, we modelled the data jointly with that of Jansen et al. (2021), whose study did not involve AI-assistance. This approach enables us to compare the DKE in our sample, where participants used AI, with the non-AI sample of Jansen et al. (2021). The model accounts for ceiling effects in performance estimates, treating scores of 20 as censored observations.
(a) Scatter-plot of Estimated LSAT Scores as a function of Achieved LSAT Scores
@@
(b) Classical Dunning-Kruger Quartile-plot. Comparison of Mean Estimated and Mean Achieved Scores Across Quartiles.
@@
Figure 5. Correlation of estimated and achieved LSAT score from different perspectives, individual 5(a) vs. quartile-level 5(b).
@@
Our model mitigates the issue around regression to the mean by explicitly modeling latent skill $\theta_{i}$ as a continuous variable with a flexible distribution. By incorporating noise $\sigma_{k}$ that scales skill level $\theta_{i}$ and bias $b_{k}$, the model allows for greater variability in judgment among low-skill participants. This scaling effect, coupled with a bias term $b_{k}$ and hierarchical priors, reduces the tendency for all participants to regress toward a single mean. For a DKE to exist, $b_{k}$ 0 and $\sigma_{k}$ 1 has to be satisfied. If only metacognitive bias is driving a DKE pattern, then $\sigma_{k}$ will be centered at 1 (values of $\sigma_{k}$ under 1 and close to zero would mean that high-performers would be less accurate in estimating their performance  [^6] To fit our data into the model, we used the STAN-sampler (Carpenter et al., 2017). Four Hamilton-Monte-Carlo chains were computed, each with 15,000 iterations and a 30% warm-up. Trace plots of the Markov-chain Monte-Carlo permutations were inspected for divergent transitions and autocorrelation, and we checked for local convergence. All Rubin-Gelman statistics (Gelman and Rubin, 1992) were well below 1.1, and the Effective Sampling Size was over 1000. We then analyzed the posterior of the model. To investigate a parameter’s distinguishability from zero, we utilized $p_{b}$, which resembles the classical $p$-value but quantifies the effect’s likelihood of being zero (for $b$) and one (for $\sigma$) or opposite (Hoijtink and van de Schoot, 2018; Shi and Yin, 2020). Effects with $p_{b}$leq 2.5% were deemed distinguishable. We also calculated the 95% High-Density Interval (HDI) for each model parameter; for visualization of prior and posterior, see Figure 6.). The bias parameter for our AI-augmented sample, $b_{AI}$, showed a median of 0.45 (95% HDI [0.32, 0.60]). The consistently positive bias indicates that individuals, when AI-augmented, tend to overestimate their abilities. We also find a bias $b_{noAI}$ for the non-AI-augmented group of Jansen et al. (2021) (Median = 0.23 (95% HDI [0.21, 0.25], $p_{b}$ = 0.0%) although when comparing posterior samples (see Figure 6), 99% of posterior samples were larger in the AI-augmented group as compared to the non-augmented sample.
@@
To understand how metacognitive noise affected self-assessment, we can turn to $\sigma_{k}$. For the non-AI-augmented group, we find a $\sigma_{noAI}$ above 1, indicating noise affecting self-assessment. This group had a median of 1.78 (95% HDI [1.69, 1.88], $p_{b}$ = 0.0%, see Figure 6B), indicating noise in judgment for the sample of Jansen et al. (2021). Combined with bias, this contributes to the DKE (see also Figure 6C for posterior predictions from the model). In comparison, our sample, which used AI to complete the task, was not affected by the DKE (Median = 1.01, 95% HDI [0.84, 1.19], $p_{b}$ = 45.66%). Given the non-overlapping distributions (0% overlap) and the small HDI’s, we can assert that with $\sigma_{AI}$ being over 1, scaling of the equation of bias and skill is not present in our sample. Hence, when augmented with AI, we observe no DKE (see again Figure 6C).
@@
Plot (a): comparison of posterior distributions for the bias parameter between the AI group and no-AI group (larger for AI-group). Plot (b): comparison of posterior distribuitions of metacognitive noise for AI and non-AI group, where the AI group shows less noise in judgement than the no-AI group. Plot (c): scatter plot showing theposterior predicted values for percent correct achieved (x-axis) and percent correct expected (y-axis) for each group. The s-shape around ideal metacognitive accuracy indicates a DKE with low-performers overestimating their performance more than high-performers.
4.5. Qualitative Data
In addition to quantitative measurements, we analyzed the qualitative data collected during the study using an inductive thematic approach (Clarke and Braun, 2017). This included both the prompts participants entered into the AI chatbot interface and their responses to an open-ended question at the end of the questionnaire. The prompts were filtered to exclude those that were a direct copy from the task, ensuring that only meaningful interactions were kept for this analysis. The remaining prompts were then inductively analyzed to identify recurring themes (section 4.5.1). Responses to the open-ended question were analyzed to explore differences in AI perception during the interactions, where recurrent themes were found (section 4.5.2).

### v2

@@ 3.1.2. Materials
Participants’ logical reasoning ability was measured with the 20 multiple-choice logical reasoning problems used by Jansen et al. (2021) to approximate the LSAT a widely recognized, real-world assessment used in high-stakes decision-making, such as law school admissions (Shultz and Zedeck, 2011; Wainer, 1995). It also serves as a benchmark in machine learning research, making it ideal for comparing AI-assisted performance (Katz et al., 2024). An example LSAT question provided to participants was: ”it has been proven that the lie detector can be fooled. If one is truly aware that one is lying, when in fact one is, then the lie detector is worthless. The author of this argument implies that: (1) The lie detector is sometimes worthless. (2) The lie detector is a useless device. (3) No one can fool the lie detector all of the time. (4) A good liar can fool the device. (5) A lie detector is often inaccurate.”.
Using the same items as Jansen et al. (2021) enabled us to compare our results to a representative sample of participants who did not use AI in the task and replicate the results of the original study by Kruger and Dunning (1999).
@@
3.1.3. Task
Participants completed 20 LSAT logical reasoning items in a randomized order. Each problem was displayed on the left-hand side of the screen, while a ChatGPT interface was presented on the right. Participants were required to interact with ChatGPT for assistance, ensuring at least one prompt per problem, before submitting their final answers and rating their confidence in their response (”How confident are you that your response is correct?”; from ”unsure” to ”certain” on a 100-step slider). Unlimited text chat interaction was enabled during the task, allowing participants to engage with ChatGPT as much as they deemed necessary. The LSAT problems were intended to assess logical reasoning abilities and did not require any prior knowledge of law to solve.
3.1.4. Apparatus
@@
We inspected the software and flow of Jansen et al. (2021) and carefully replicated it, integrating a side-by-side view of ChatGPT and survey interface (Figure 2). We used ChatGPT-4o (gpt-4o-2024-05-13) due to its widespread use in enhancing cognitive task performance (Bastani et al., 2024; Draxler et al., 2024, 2023). A custom interface (Figure 2) was built to log user interactions, enabling qualitative chat analysis. The application automatically collected survey responses and chat logs and recorded them at the end of the study. We included a button to copy each problem and its answer options to the clipboard so they could be easily pasted into the chat.
3.1.5. Procedure
Upon entering the study on Prolific, participants were redirected to our application. After consenting to participate, we asked our participants to complete the user expectations questionnaire. Then, they were briefly introduced to the task and allowed to test the chat interaction. Afterwards, participants engaged in a task to assess their logical reasoning skills by solving a series of LSAT problems. Before submitting their final answers, they were asked to interact with ChatGPT if assigned to the AI group, ensuring they provided at least one prompt per problem. After completing each question, participants rated their confidence in their response using a 100-step slider ranging from “unsure” to “certain.” In the AI group, participants were permitted unlimited text-based interaction with ChatGPT, allowing participants to seek as much assistance as they felt necessary during the problem-solving process. After solving the problems in a randomized order (3.1.3), participants were again asked to complete the expectations questionnaire in the past tense. They also responded to the SNAIL questionnaire (Laupichler et al., 2023) and filled in their demographic information, including age, gender, occupation, education level, English proficiency, and whether they had taken the LSAT before. Study 1 took, on average, 42 minutes to complete.
3.1.6. Analysis
@@
We analyzed the data in five steps. First, we compared our sample performance to that of Jansen et al. (2021) and to the performance of AI alone. This allowed us to analyze where AI augments human performance (i.e. human-AI interaction outperforms a no-AI group) and human-AI synergy (i.e. human-AI interaction outperforms AI). We analyze both average performance and its distribution. Second, we analyzed metacognitive accuracy on a task level – the difference between objective and estimated performance [^1] In line with Ehrlinger et al. (2008), we focus on numeric estimates of performance after the task, not relative performance in comparison to others. comparing human-AI interaction to a no-AI group – and trial-level confidence ratings to assess metacognitive sensitivity. With this, we can analyze metacognitive performance of users when interacting with AI globally (accuracy) after the tasks and locally (sensitivity) after each decision. Note that for Jansen et al. (2021), the question was (”How many of the 20 logical reasoning problems do you think you solved correctly?”) while in our sample using AI, we asked: ”Using the AI, how many of the 20 logical reasoning problems do you think you solved correctly?”. Next, we correlated metacognitive performance metrics with performance and AI literacy measures to explore what predicts low metacognitive performance in human-AI interaction (for a similar analysis approach see McIntosh et al. (2019)). Fourth, we used a computational model of performance and performance assessments to compare our sample and Jansen et al. (2021) to estimate how AI affects the DKE. Lastly, we qualitatively analyzed participant strategies and how they conceptualize the human-AI relation [^2] Given the large samples in our study and the ceiling effects encountered 5(a), we do not test for the normality of residuals of our variables. Instead, we model the data using a final Bayesian computational framework, which allows for more flexible assumptions and can account for ceiling effects in our main analysis. This approach provides more robust estimates by incorporating uncertainty in a probabilistic manner rather than relying on strict parametric assumptions.. We use frequentist statistics ($\alpha$= 5%) for simple statistical tests (e.g., paired and unpaired t-test and Pearson correlation) in the first four analyses and Bayesian statistical modelling for the computational model. Note that we focus only on these analyses in the exploration of our data for the current paper; however, we encourage the re-analysis and further exploration in the openly available dataset https://osf.io/svax9/?view_only=cd176c08c2d4422baabf6bf815018bd7.
3.2. Results
@@
Note: M (SD) and the sample size (n). Scale for the assessment of non-experts’ AI literacy (SNAIL). According to task context, rank % instructions are scaled as follows: marking 90% means you perform better than 90% of participants, marking 10% means you perform better than only 10% of participants, and marking 50% means you will perform better than half of the participants.- indicates no variation, e.g., when all participants in a quantile had the same value.
3.2.1. Human-AI Composite Performance
Figure 3. Comparison of performance scores between participants interacting with ChatGPT and a dataset without AI (Jansen et al., 2021). The blue density curve represents the distribution of performance scores in a sample of 246 participants using ChatGPT, showing a peak in performance at around 14 points. The yellow density curve corresponds to a larger sample (n = 3543) from the Jansen et al. (2021) dataset, with lower overall performance scores. The vertical dashed line indicates the mean performance score in the ChatGPT simulation.
@@ 3.2.1. Human-AI Composite Performance
The graph compares performance scores between participants using ChatGPT and those without AI assistance from the Jansen et al. dataset using density plots. The ChatGPT group shows higher performance, peaking around 14 points, while the no-AI group has lower overall scores. A dashed line marks the mean performance in the ChatGPT simulation. One can clearly see that performance is increased for ChatGPT interaction and that this distribution is much more narrow than Jansen’s data.
@@
To see whether there is a synergy effect of using ChatGPT in the LSAT (i.e. Human-AI performance ¿ AI performance) we compared the average ChatGPT performance (100 runs at $M$ = 13.65;) to our user’s average performance ($M$ = 12.98, $SD$ = 2.88). We find a significant difference, with participants performing slightly worse than ChatGPT alone $t(245)=-3.66,p<.001,d=-0.23$ in the task. Next, we compared our sample’s performance to Jansen et al. (2021) representative sample of 3543 participants, who completed the same task without any assistance ($M$ = 9.45, $SD$ = 3.59). We find that in our sample, participants performed significantly better with ChatGPT assistance $t(245)=19.23,p<.001,d=1.23$ as compared to the Jansen et al. (2021) sample. Therefore, while, on average, there is no human-AI synergy, we do find that ChatGPT use can augment human performance for solving the LSAT logical problems (i.e. human-AI performance ¿ human performance).
@@
Looking at individual performance, we can find indications of human-AI synergy. The difference between our sample’s and ChatGPT’s performance is rather small, at less than one point. 55.28% (136 of 246) of our participants performed better than ChatGPT. However, 89.43% (220 of 246) in our sample performed better than the average score of the Jansen et al. (2021) sample (see also Figure 3).
@@
Therefore, while overall performance increased with the use of ChatGPT augmenting the human ability to solve LSAT problems, on average, we do not find a human-AI synergy. The composite performance of ChatGPT and the participant overtook the performance of ChatGPT alone for only slightly more than half of the participants in our sample. With our prototype’s ability to enhance performance established in human augmentation but not synergy, we can now focus on investigating metacognitive abilities.
3.2.2. Metacognitive accuracy and sensitivity
Participants were inaccurate in assessing their performance after task completion, as indicated in the item “Using the AI, how many of the 20 logical reasoning problems do you think you solved correctly?”, see also Table 1. On average, they estimated solving about 17 out of 20 items ($M$ = 16.50, $SD$ = 3.72). This overestimation of about 4 points could be distinguished from 0, $t(256)=14.98,p<.001,d=0.93$.
@@ 3.2.2. Metacognitive accuracy and sensitivity
To see whether participants track information in each trial, we turn to metacognitive sensitivity that we estimate from confidence ratings when making the decision. Mean confidence (rated on a scale from 0-100) for correct answers was 82.49 (14.24) and for incorrect answers 77.00 (16.52), t(244) = 8.21, p ¡ .001, d = 0.52. To evaluate sensitivity more granularly, we conducted a Receiver Operating Characteristic (ROC) analysis. ROC analysis is a technique used to assess the performance of a judgment by plotting the true positive rate against the false positive rate across different thresholds. Here, we applied ROC analysis to understand how well participants’ confidence scores predicted whether their responses were correct (see Figure 4A).
By using ROC analysis, we obtain a metric, the area under the curve (AUC) that quantifies how effectively a participant’s confidence ratings differentiate between correct and incorrect responses. An AUC value of 0.5 indicates no better-than-chance discrimination, while higher values reflect greater sensitivity—meaning the participant’s confidence reliably tracks correctness. Thus, the ROC analysis provides a nuanced individual, trial-level measure of metacognitive accuracy beyond simple average confidence or aggregate performance estimates.
The mean AUC was .62 ($SD$ = 11.2) which could be distinguished from 0.5 ($t(244)=16.02,p<.001,d=1.02$). While most participants’ (210 out of 246; 85.37%) metacognitive AUC values are above .50 (random guessing). This means that confidence scores indicate participants’ metacognitive sensitivity on a trial level in human-AI interaction. Note that for the remaining 36 participants’ confidence ratings were not able to distinguish between correct or incorrect trials (for the distribution of AUC values, refer to Figure 4B). For these participants, confidence judgments were effectively random or worse than random chance, indicating that they tended to be as confident or even more confident about incorrect responses than correct ones. This pattern suggests a miscalibration in their metacognitive judgments, where confidence fails to serve as a reliable indicator of actual performance. Thus, our sample exhibits very low metacognitive sensitivity and, in consequence, low metacognitive monitoring.
@@
Two plots related to metacognitive sensitivity. Plot A displays Receiver Operating Characteristic (ROC) curves for participants, illustrating the relationship between True Positive Rate (Sensitivity) and False Positive Rate. With most participants performing better than chance. The dashed diagonal line represents random guessing. For each participant, we generated a ROC curve, as well as one from pooled responses, which illustrates the trade-off between the true positive rate. Plot B shows the distribution of Area Under the Curve (AUC) values across all participants, with a peak around 0.6, indicating variability in metacognitive sensitivity. Most participants perform slightly above chance level (AUC = 0.5).
3.2.3. Correlation of metacognitive ability, performance, and AI literacy (SNAIL)
A number of significant relationships were found when correlating several metacognitive indices with LSAT performance and AI literacy, see Table 2. We observed a positive relation between performance and participants’ average confidence estimates. Participants who perform well are also more confident on average. However, those who are, on average, more confident also overestimate their performance due to increased metacognitive bias. This is probably due to the relationship between SNAIL factors and performance estimates, where participants who express more technical knowledge and more critical appraisal also estimate their performance to be relatively higher. However, those with high technical understanding are also less accurate in their metacognitive judgements. All SNAIL factors correlated positively to average confidence. Note that these correlations are rather small and should thus be interpreted with caution. Metacognitive sensitivity (AUC and $\Delta conf$) was not related to AI literacy, performance or metacognitive accuracy.
@@ 3.2.3. Correlation of metacognitive ability, performance, and AI literacy (SNAIL)
$\Delta conf$0.04 -0.03 0.01
$\mu conf$ 0.24*** 0.46*** 0.27*** -0.10
AUC 0.03 0.05 0.03 0.59*** -0.08
SNAIL TU 0.21** 0.17** -0.06 -0.12 0.13* -0.10
SNAIL CA 0.10 0.14* 0.06 0.04 0.25*** 0.03 0.49***
SNAIL PA 0.05 0.10 0.06 0.01 0.24*** 0.05 0.57*** 0.81***
@@
3.2.4. AI use cancels the Dunning-Kruger effect
@@
We calculated the difference between quantiles to test whether metacognitive accuracy was worse in the low-scoring quantile than in the high-scoring quantile. Both quantile’s metacognitive accuracy differed from 0, (Q1: $t(109)=-10.15,p<.001,d=-0.97$, Q4: $t(21)=-3.64,p=.002,d=-0.78$), probably due to the overall bias. However, we found that the difference for Q1 is larger than for Q4, $t(130)=2.79,p=.006,d=0.49$ (see also 5(b) and Table 1). Note that this pattern of effect could be driven by metacognitive bias alone. To establish a DKE, we must first quantify the metacognitive noise in our sample. To do so, we employ a Bayesian computational model  [^4] For a guide on Bayesian techniques, see (Schad et al., 2021; van de Schoot et al., 2021; Dix, 2022; Bürkner, 2017; Kay et al., 2016), we used the tutorial of Nathaniel Haine’s as a starting point for our modelling efforts: http://haines-lab.com/post/2021-01-10-modeling-classic-effects-dunning-kruger/. a hierarchical Bayesian model to jointly estimate participants’ objective and perceived performance while accounting for latent skill, metacognitive bias, and metacognitive noise  [^5] For a theoretical model, see Burson et al. (2006). To allow for a baseline comparison of the DKE, we modelled the data jointly with that of Jansen et al. (2021), whose study did not involve AI. This approach enables us to compare the DKE in our sample, where participants used AI, with the non-AI sample of Jansen et al. (2021). The model accounts for ceiling effects in performance estimates, treating scores of 20 as censored observations.
(a) Scatter-plot of Estimated LSAT Scores as a function of Achieved LSAT Scores for the AI group in Study 1.
@@
(b) Classical Dunning-Kruger Quartile-plot. Comparison of Mean Estimated and Mean Achieved Scores Across Quartiles for the AI group in Study 1.
@@
Figure 5. Correlation of estimated and achieved LSAT score from different perspectives, individual 5(a) vs. quartile-level 5(b) for the AI group in Study 1.
@@
Plot (a): scatter plot of estimated LSAT scores as a function of achieved LSAT scores across quartiles (Q1 to Q4). While some participants were accurate in their estimates, some participants were very wrong in estimating their performance accurately. Plot (b): quartile-level Dunning-Kruger plot, showing that estimated scores are consistently higher than achieved scores across all quartiles with a spread for low performers and closer match of estimates and performance in high performers.
@@
Our model mitigates the issue around regression to the mean by explicitly modelling latent skill $\theta_{i}$ as a continuous variable with a flexible distribution. By incorporating noise $\sigma_{k}$ that scales skill level $\theta_{i}$ and bias $b_{k}$, the model allows for greater variability in judgment among low-skill participants. This scaling effect, coupled with a bias term $b_{k}$ and hierarchical priors, reduces the tendency for all participants to regress toward a single mean. For a DKE to exist, bias that is $b_{k}$ 0 and noise that is $\sigma_{k}$ 1 has to be satisfied. If only metacognitive bias is driving a DKE pattern, then $\sigma_{k}$ will be centered at 1 (values of $\sigma_{k}$ under 1 and close to zero would mean that high-performers would be less accurate in estimating their performance  [^6] To fit our data into the model, we used the STAN-sampler (Carpenter et al., 2017). Four Hamilton-Monte-Carlo chains were computed, each with 15,000 iterations and a 30% warm-up. Trace plots of the Markov-chain Monte-Carlo permutations were inspected for divergent transitions and autocorrelation, and we checked for local convergence. All Rubin-Gelman statistics (Gelman and Rubin, 1992) were well below 1.1, and the Effective Sampling Size was over 1000. We then analyzed the posterior of the model. To investigate a parameter’s distinguishability from zero, we utilized $p_{b}$, which resembles the classical $p$-value but quantifies the effect’s likelihood of being zero (for $b$) and one (for $\sigma$) or opposite (Hoijtink and van de Schoot, 2018; Shi and Yin, 2020). Effects with $p_{b}$leq 2.5% were deemed distinguishable. We also calculated the 95% High-Density Interval (HDI) for each model parameter; for visualization of prior and posterior, see Figure 6.). The bias parameter for our AI-interaction sample, $b_{AI}$, showed a median of 0.45 (95% HDI [0.32, 0.60]). The consistently positive bias indicates that individuals, when using AI, tend to overestimate their abilities. We also find a bias $b_{noAI}$ for the non-AI group of Jansen et al. (2021) (Median = 0.23 (95% HDI [0.21, 0.25], $p_{b}$ = 0.0%) although when comparing posterior samples (see Figure 6), 99% of posterior samples were larger in the AI group as compared to the non-augmented sample.
@@
To understand how metacognitive noise affected self-assessment, we can turn to $\sigma_{k}$. For the non-AI group, we find a $\sigma_{noAI}$ above 1, indicating noise affecting self-assessment. This group had a median of 1.78 (95% HDI [1.69, 1.88], $p_{b}$ = 0.0%, see Figure 6B), indicating noise in judgment for the sample of Jansen et al. (2021). Combined with bias, this contributes to the DKE (see also Figure 6C for posterior predictions from the model). In comparison, our sample, which used AI to complete the task, was not affected by the DKE (Median = 1.01, 95% HDI [0.84, 1.19], $p_{b}$ = 45.66%). Given the non-overlapping distributions (0% overlap) and the small HDI’s, we can assert that with $\sigma_{AI}$ being over 1, scaling of the equation of bias and skill is not present in our sample. Hence, when augmented with AI, we observe no DKE (see again Figure 6C).
@@
Plot (a): comparison of posterior distributions for the bias parameter between the AI group and no-AI group (larger for the AI group). Plot (b): comparison of posterior distributions of metacognitive noise for AI and non-AI groups, where the AI group shows less noise in judgement than the no-AI group. Plot (c): scatter plot showing the posterior predicted values for percent correct achieved (x-axis) and percent correct expected (y-axis) for each group. The s-shape around ideal metacognitive accuracy indicates a DKE with low-performers overestimating their performance more than high-performers.
3.2.5. Qualitative Data
In addition to quantitative measurements, we analyzed the qualitative data collected during Study 1 using an inductive thematic approach (Clarke and Braun, 2017). This included both the prompts participants entered into the AI chatbot interface and their responses to an open-ended question at the end of the questionnaire. The prompts were filtered to exclude those that were a direct copy from the task, ensuring that only meaningful interactions were kept for this analysis. The remaining prompts were then inductively analyzed to identify recurring themes (section 3.2.6). Responses to the open-ended question were analyzed to explore differences in AI perception during the interactions, where recurrent themes were found (section 3.2.6).